# ViViDex: Learning Vision-based Dexterous Manipulation from Human Videos

### v1

@@ Abstract
In this work, we aim to learn a unified vision-based policy for a multi-fingered robot hand to manipulate different objects in diverse poses. Though prior work has demonstrated that human videos can benefit policy learning, performance improvement has been limited by physically implausible trajectories extracted from videos. Moreover, reliance on privileged object information such as ground-truth object states further limits the applicability in realistic scenarios. To address these limitations, we propose a new framework ViViDex to improve vision-based policy learning from human videos. It first uses reinforcement learning with trajectory guided rewards to train state-based policies for each video, obtaining both visually natural and physically plausible trajectories from the video. We then rollout successful episodes from state-based policies and train a unified visual policy without using any privileged information. A coordinate transformation method is proposed to significantly boost the performance. We evaluate our method on three dexterous manipulation tasks and demonstrate a large improvement over state-of-the-art algorithms. The project page is available at https://zerchen.github.io/projects/vividex.html.
Figure 1: The overall framework of our method for learning dexterous manipulation skills from human videos. It consists of three modules: the extraction of the reference trajectory from human videos, the trajectory-guided state-based policy learning using RL, and the vision-based policy learning using behavior cloning.
@@ I Introduction
Humans possess a remarkable ability to manipulate diverse objects using their hands based on visual perception. Nevertheless, compared with parallel grippers [1, 2, 3, 4, 5, 6], achieving comparable levels of dexterous manipulation with multi-fingered robot hands remains an ongoing challenge [7]. It requires to convert visual signals to high-dimensional control commands.
@@
In recent years, the advancement of deep learning (DL) [8] and reinforcement learning (RL) [9, 10] has led to significant progress in learning-based algorithms for dexterous manipulation such as in-hand rotation [11, 12, 13], solving Rubik’s cube [14] and playing the piano like humans [15, 16]. However, it proves challenging to train RL policies as RL training heavily relies on intricate reward engineering [10, 17] and extensive computational resources [14]. Furthermore, unnatural behaviors with high rewards can emerge from RL training [18]. To address the limitations of RL training, some previous works [19, 20, 21, 22] have turned to imitation learning using robot demonstrations collected by teleoperation. While these approaches enhance training efficiency, they demand substantial human efforts to collect diverse robot demonstration data and are, hence, difficult to scale.
@@
As human videos capturing hands manipulating objects are more abundant and easy to acquire, an increasing research attention has been paid to leveraging human videos for learning dexterous manipulation skills [23, 24, 25, 26, 27]. For example, DexMV [23] proposes to automatically extract robot and object poses from human videos, and then employs the data to accelerate RL via demonstration augmented policy gradient (DAPG) [19]. However, since the extracted robot and object poses are noisy, the method needs hundreds of human videos to learn the manipulation of a single object and still requires reward engineering for different tasks. Additionally, existing works [23, 25, 28] take advantage of privileged information of objects in policy learning such as known object CAD models and ground-truth object poses, which are non-trivial to obtain from raw visual sensors in real-world scenarios.
@@
In this work, we propose a new framework called ViViDex to leverage human Videos for Vision-based Dexterous manipulation policy learning. It consists of three modules as illustrated in Figure 1: reference trajectory extraction, trajectory guided state-based policy learning, and unified vision-based policy learning. We first obtain reference trajectories for hands and objects from human videos following prior work [23]. Though such trajectories are noisy and not directly suitable for robot control, they provide examples of natural hand-object interactions that we use as guidance. Our second module refines and recovers physics properties of the reference trajectory. We use RL for this purpose by training a state-based policy for each reference trajectory. We design a novel reward to keep pose similarity in the reference trajectory while solving the task. We also augment the trajectory during training to generalize to a broader range of object poses beyond the pose in the reference trajectory. Finally, we rollout episodes from optimized state-based policies and distill successful episodes to a unified vision-based policy without any privileged object information. A coordinate transformation method is proposed to improve the vision-based policy, which transforms the input 3D point clouds into hand joints and the target coordinate systems to capture fine-grained interaction features and increase awareness of the target position. We validate the effectiveness of our system on three dexterous manipulation tasks relocation, pouring and placing inside. Our ViViDex approach significantly outperforms the state-of-the-art method DexMV [23] while using significantly less human demonstrations.
@@
$\bullet$ Experiments on three dexterous manipulation tasks demonstrate the effectiveness of the ViViDex approach.
@@ II Related Work
Dexterous manipulation. The multi-fingered robotic hands empower robots to perform delicate operations on objects like humans. Previous works can be generally grouped into two categories: trajectory optimization methods [29, 30, 31, 32, 33] and data-driven learning methods [19, 23, 24, 28, 34, 35, 36, 37]. Trajectory optimization methods usually require well-defined dynamic models for the robot and the manipulated object, which are usually not easy to obtain in practice. Data-driven methods instead train neural network policies for manipulation, allowing to learn generalizable skills. Most of the works [19, 28, 35, 38, 39, 40] use reinforcement learning (RL) algorithms to train policies. Nagabandi et al. [39] and Hansen et al. [40] propose to improve the sample efficiency in RL by jointly learning the dynamic model and the control policy. To accelerate the convergence of RL, Rajeswaran et al. [19] introduce the demo augmented policy gradient algorithm (DAPG) and train RL with expert demonstrations. Instead of manually collecting expert demonstrations, ILAD [28] derives demonstration data via an affordance model [41] and trajectory optimization [42]. Xu et al. [35] and Wan et al. [38] decompose the dexterous grasping into two sub-tasks: the grasp proposal generation and the goal-conditioned policy training and largely improve the generalization of the learned policy. However, RL is data-inefficient and requires well-designed rewards, which often results in unrealistic object manipulation actions. To address these problems, more recent works [23, 25, 26, 27] attempt to learn policies using human demonstrations and combine imitation learning with RL.
@@
Learning robotic manipulation from human videos. Human videos naturally provide rich human hands interactions with diverse objects [43, 44, 45, 46, 47]. Equipping robots with the capacity to acquire manipulation skills by simply watching human videos has been an attractive direction [48, 49, 50, 51, 52]. Some works [34, 53, 54, 55] propose to use human videos to pre-train a visual representation for the policy learning. R3M [34] proposes to pre-train CNNs with egocentric human videos [56] and perform the imitation learning based on the learned visual representations. Xiao et al. [53] and Radosavovic et al. [54] propose to leverage in-the-wild human videos to pre-train the vision transformer and demonstrate the improvement on different manipulation tasks. The recent work H-InDex [55] employs 3D human pose estimation from human videos as a proxy task to pre-train the visual representation for dexterous manipulation tasks. Another line of works [57, 58, 59, 60] focus on learning the reward function from human videos. Chen et al. [57] introduce DVD, a domain-agnostic video discriminator, to learn multi-task reward functions. Alakuijala et al. [58] propose a task-agnostic reward function by using unlabeled human videos. Chane-Sane et al. [59] formulate the reward learning as a retrieval task and propose a data-driven approach to map human demonstrations to unseen robotic manipulation tasks. Xiong et al. [60] propose to translate human demonstration videos into robot demonstration videos and use the keypoint-based representation to define the distance metric for the policy execution. Some other works [23, 24, 25, 26, 61] explore how to explicitly leverage human videos to facilitate learning dexterous manipulation skills. Mandikal et al. [61, 24] propose to utilize object affordances priors and human grasping priors to improve the dexterous grasping policy. Shaw et al. [26] propose to first pre-train the policy with visual, action and physical priors learned from human videos and finetune the policy with teleoperated demonstrations to make it generalize to different robotic tasks. DexMV [23] and DexRepNet [25] extracts demonstrations data from human videos and uses DAPG algorithm [19] to launch the RL training with demonstrations. In this work, not limited to the simple dexterous grasping action [24, 25, 35, 38, 61], we aim to solve more complex tasks by utilizing human videos. To additionally address limitations of previous works [23, 25, 28] that require object states (e.g., object poses) during the policy execution, we propose learning vision-based policies that achieve high manipulation accuracy while only using proprioceptive robot information and depth observations in the form of point clouds.
@@ III Method
TABLE I: The state definition. The left part summarizes robot states. The right part defines object states.
Parameters Descriptions Parameters Descriptions
$\bm{q}_{r}\in\mathbb{R}^{30}$ robot motor positions $\bm{q}_{o}\in\mathbb{R}^{6}$ object canonical positions
$\bm{\dot{q}}_{r}\in\mathbb{R}^{30}$ robot motor velocities $\bm{\dot{q}}_{o}\in\mathbb{R}^{6}$ object canonical velocities
$\bm{x}_{\text{r}}\in\mathbb{R}^{16\times 3}$ robot joint positions $\bm{x}_{o}\in\mathbb{R}^{3}$ object center position
$\bm{\theta}_{\text{r}}\in\mathbb{R}^{16\times 4}$ robot joint rotations $\bm{\theta}_{o}\in\mathbb{R}^{4}$ object global rotation
$\bm{\dot{x}}_{\text{r}}\in\mathbb{R}^{16\times 3}$ robot linear velocities $\bm{\dot{x}}_{\text{o}}\in\mathbb{R}^{3}$ object linear velocity
$\bm{\omega}_{\text{r}}\in\mathbb{R}^{16\times 3}$ robot angular velocities $\bm{\omega}_{\text{o}}\in\mathbb{R}^{3}$ object angular velocity
$\bm{t}_{\text{r}}\in\mathbb{R}^{7\times 3}$ robot target translations $\bm{t}_{\text{o}}\in\mathbb{R}^{3}$ object target translations
This work aims to train vision-based dexterous manipulation policies. However, jointly learning the visual representation and the control policy is challenging. Therefore, we first train the state-based policy and then distill the generated data to a vision-based policy. To train the state-based policy, as the pure RL training is ineffective [23, 25, 28], we extract reference trajectories from human videos and propose trajectory-guided RL to improve the performance. In the following, we first present details about the reference trajectory generation in Section III-A. Then, we introduce our state-based policy learning algorithm in Section III-B. Finally, we describe how we distill multiple state-based policies into the visual policy and its model architecture in Section III-C.
@@ III-A Reference Trajectory Extraction
In this work, we use the DexYCB [43] video dataset, which captures human hand grasping of objects. We first extract hand and object poses from the video. The hand poses and shapes are represented by MANO [62] and result in 3D locations of joints $\bm{\psi}_{h}\in\mathbb{R}^{21\times 3}$. Then, we aim to retarget the human hand pose to the robot hand pose. We use the Adroit anthropomorphic robot [19], which consists of 30 degrees of freedom $\bm{q}_{r}\in\mathbb{R}^{30}$. Following [23, 63, 64], we formulate the hand motion retargeting from a video of length $T$ as an optimization problem and define its objective function as:
$\begin{split}&\min_{\bm{q}_{r}^{t}}\sum_{t=1}^{T}\Big{\|}\bm{\hat{x}}_{rt}^{t}(\bm{q}_{r}^{t})-\bm{{\psi}}_{ht}^{t}\Big{\|}^{2}_{2}\\ &+\Big{\|}\bm{\hat{x}}_{rp}^{t}(\bm{q}_{r}^{t})-\bm{{\psi}}_{hp}^{t}\Big{\|}^{2}_{2}+\alpha\Big{\|}\bm{q}_{r}^{t}-\bm{q}_{r}^{t-1}\Big{\|}^{2}_{2},\end{split}$ (1)
where $\bm{{\psi}}_{ht}^{t}\in\mathbb{R}^{5\times 3}$ and $\bm{{\psi}}_{hp}^{t}\in\mathbb{R}^{5\times 3}$ denote human hand tip and middle phalanx positions, respectively. We solve their counterparts $\bm{\hat{x}}_{rt}^{t}$ and $\bm{\hat{x}}_{rp}^{t}$ in the robot via forward kinematics according to given ${\bm{q}_{r}^{t}}$. In the first two terms, we aim to minimize the difference between the 3D locations of robot joints to those in the human hand. We also use a regularization term to avoid sudden changes in the robot pose, where we empirically set $\alpha$ and ${\bm{q}_{r}^{0}}$ to 4e-3 and the mean pose of its motion limits, respectively. We use the NLopt solver [65] for the optimization and port robot and object motions into the simulator to generate reference trajectories, which are only visually plausible but not physically plausible.
Figure 2: The illustration of different stages for the reference trajectory. The pre-grasp stage (a) describes how the robot approaches the object. The manipulation stage (b) focuses on how the robot hand manipulates the object.
@@ III-B State-based Policy
We employ the generated reference trajectory as the reward function for RL to constrain the robot hand and object motions and train a state-based policy to recover physically plausible trajectories. The network architecture for the state-based policy consists of actor and critic MLPs [66]. It takes both robot and object states described in Table I as inputs and predicts the robot control commands. Different from DexMV [23] that needs different reward functions for different manipulation tasks, our method can solve different tasks with the same reward functions. In the following, we will describe our reward functions and the trajectory augmentation for RL.
Reward functions. To define the reward function, we divide the reference trajectory into two stages shown in Figure 2: the pre-grasp stage and the manipulation stage. The recent work PGDM [27] focuses on the manipulation stage and initializes the robot hand at a given pre-grasp configuration for RL, which is a simplified setting compared with ours.
During the pre-grasp stage, the robot hand approaches the object without making physical contacts. During RL training, we require that the robot approaches the object in a similar way as humans and define the following reward function:
@@
where $T_{p}$ is the length of pre-grasp steps, and $\bm{\hat{x}}_{rt}^{t}\in\mathbb{R}^{5\times 3}$ denotes the robot finger tip positions at the timestep $t$ in the reference trajectory $\bm{{x}}_{rt}^{t}$ is the current robot tip locations.
@@
When the robot successfully reaches its pre-grasp configuration, the episode starts its manipulation stage. In this stage, the robot manipulates the object and aims to bring it to the desired target configuration. Here, we define the reward function to constrain the robot and object motions jointly:
$\begin{split}&R_{m}=\sum_{t=T_{p}+1}^{T_{r}}\lambda_{1}\cdot{\rm exp}(-\alpha_{1}\cdot\Big{\|}\bm{{x}}_{rt}^{t}(\bm{q}_{r}^{t})-\bm{\hat{x}}_{rt}^{t}\Big{\|}^{2}_{2})\\ &+\lambda_{2}\cdot{\rm exp}(-\alpha_{2}\cdot(\Big{\|}\bm{x}_{o}^{t}-\bm{\hat{x}}_{o}^{t}\Big{\|}^{2}_{2}+\alpha_{3}\cdot\phi(\bm{\theta}_{o}^{t},~{}\bm{\hat{\theta}}_{o}^{t})))\\ &+\sum_{i=1}^{5}\lambda_{3}\cdot\mathds{1}^{\rm contact}_{\rm finger~{}i}+\lambda_{4}\cdot\mathds{1}^{\rm lift}_{\rm object},\end{split}$ (3)
@@
where $T_{r}$ refers to the length of the whole reference trajectory. The first term of (3) constrains the hand motions similar to (2). The second term constrains the object motions, and $\bm{x}_{o}^{t}$ and $\bm{\hat{x}}_{o}^{t}$ represent the current object position and the reference object position at timestep $t$. $\phi(\cdot)$ computes the angular distance between the current object orientation with its reference $\bm{\hat{\theta}}_{o}^{t}$. In the third term, we encourage contacts between each finger of the robot and the object. The forth term assigns bonus points when the object is lifted off the table. The second and forth terms are introduced in [27], and we use the same hyper-parameters (i.e., $\lambda_{2}=10$, $\alpha_{2}=50$, $\alpha_{3}=0.1$ and $\lambda_{4}=2.5$). For our proposed first and third terms, we empirically set $\lambda_{1}$, $\alpha_{1}$ and $\lambda_{3}$ to 4, 10 and 0.5, respectively.
When the robot and object reaches their targets, we want them to remain stable in their target configurations for a while, which is especially important for the pour task to make sure that all particles in the mug have enough time to drop into the target container. To achieve this when $t~{}>~{}T_{r}$, we still use (3) and constrain the robot and object to stay in their final target configurations until the end of the episode.
Figure 3: Illustration of our vision-based policy architecture.
@@
Reference trajectory augmentation. By using our proposed reward functions, though the state-based policy can successfully imitate the hand and object motions from a video, it still cannot generalize to different initial object positions, rotations and target positions. To make our policy applicable to different initial object configurations and target positions, we introduce our reference trajectory augmentation strategies during RL training. To deal with different initial object positions or rotations, we could accordingly transform the whole reference trajectory. When we randomly set the target object position, we augment the original object trajectory through interpolations aimed at reaching the specified target position. For pour and place inside tasks, we first design the desired final object configuration. In the pour task, we ensure that the mug is tilted appropriately, which allows the particles inside to effortlessly drop into the target container. Our goal for place inside is to accurately position the banana within the container. Then, we interpolate the object position and orientation to reach the desired configuration. Similarly, we can also interpolate hand motions and use our proposed reward functions to optimize the policy.

### v2

@@ Abstract
In this work, we aim to learn a unified vision-based policy for multi-fingered robot hands to manipulate a variety of objects in diverse poses. Though prior work has shown benefits of using human videos for policy learning, performance gains have been limited by the noise in estimated trajectories. Moreover, reliance on privileged object information such as ground-truth object states further limits the applicability in realistic scenarios. To address these limitations, we propose a new framework ViViDex to improve vision-based policy learning from human videos. It first uses reinforcement learning with trajectory guided rewards to train state-based policies for each video, obtaining both visually natural and physically plausible trajectories from the video. We then rollout successful episodes from state-based policies and train a unified visual policy without using any privileged information. We propose coordinate transformation to further enhance the visual point cloud representation, and compare behavior cloning and diffusion policy for the visual policy training. Experiments both in simulation and on the real robot demonstrate that ViViDex outperforms state-of-the-art approaches on three dexterous manipulation tasks. Qualitative results are available at: zerchen.github.io/projects/vividex.html.
Figure 1: The overall framework of our method for learning dexterous manipulation skills from human videos. It consists of three steps: extraction of reference trajectories from human videos, trajectory-guided state-based policy learning using RL, and vision-based policy learning using either the behavior cloning or the 3D diffusion policy.
@@ I Introduction
People possess a remarkable ability to manipulate objects effortlessly with their hands, guided by visual perception. Despite significant progress, replicating the dexterity of human hands with multi-fingered robot hands remains challenging [1, 2, 3, 4, 5]. This is largely due to the complexity of translating visual signals into the high-dimensional control commands required for dexterous manipulation.
@@
Recent developments in deep learning (DL) [6] and reinforcement learning (RL) [7, 8] have enabled significant progress in learning-based algorithms for dexterous manipulation such as in-hand rotation [9, 10, 11], solving Rubik’s cube [12] and playing the piano [13, 14]. However, it proves challenging to train RL policies as RL training heavily relies on intricate reward engineering [8] and extensive computational resources [12]. Furthermore, unnatural behaviors with high rewards can emerge from RL training [15]. To address the limitations of RL, some previous works [16, 17, 18, 19, 20, 21] have turned to imitation learning using robot demonstrations collected by teleoperation. While these approaches enhance training efficiency, they demand substantial human efforts to collect diverse robot demonstrations and are, hence, difficult to scale.
@@
As human videos capturing hands manipulating objects [22, 23] are abundant and easy to acquire, there is a growing trend in robotics research to utilize human video demonstrations for learning dexterous manipulation skills [24, 25, 26, 27, 28]. For example, DexMV [24] proposes to extract robot and object poses from human videos, and then employs the data to accelerate RL training via demonstration augmented policy gradient [16]. However, since the extracted robot and object poses are noisy, the method needs hundreds of human videos to learn the manipulation of a single object and requires reward engineering for different tasks. Moreover, existing methods [24, 26, 29] often leverage privileged information of objects in policy learning such as ground-truth object CAD models and object poses, which are non-trivial to obtain from visual sensor data in real-world scenarios.
@@
In this work, we propose ViViDex, a new framework making use of human Videos for Vision-based Dexterous manipulation. ViViDex consists of three modules as illustrated in Figure 1. We first obtain human hand and object trajectories from video demonstrations [24]. Though such trajectories are noisy and not directly usable for robot control, they provide examples of natural hand-object interactions. Our second module then refines the reference trajectories to be physically plausible. We train a state-based policy with RL for each reference trajectory. A novel trajectory-guided reward is proposed to keep pose similarity in the reference trajectory while solving the task. We also augment trajectories during training to generalize to a broader range of object poses beyond the pose in the reference trajectory. Finally, we rollout episodes from optimized state-based policies and distill successful episodes to a unified vision-based policy using no privileged object information. To this end, we propose to improve visual representation by transforming input 3D point clouds into hand-centered coordinate systems. We evaluate ViViDex on three dexterous manipulation tasks relocation, pour and placing inside in simulation and demonstrate significant improvements over the state-of-the-art DexMV [24] while using significantly less human demonstrations. Our visual policy also achieves good performance both in simulation and on a real robot for both seen and unseen objects.
@@
$\bullet$ Extensive simulation and real robot experiments demonstrate the effectiveness of our proposed ViViDex approach.
@@ II Related Work
Dexterous manipulation. Multi-fingered robotic hands enable robots to perform delicate manipulation operations on objects. Previous work addressed dexterous manipulation by trajectory optimization [30, 31, 32, 33, 34, 35, 36] and data-driven learning methods [16, 24, 25, 29, 37, 38, 39, 40, 41, 42, 43]. Trajectory optimization methods usually require well-defined dynamic models for the robot and the manipulated object, which are sometimes hard to obtain in practice. Data-driven methods instead directly train neural policies given robot data. Most of the works [44, 45, 16, 29, 38, 46, 47, 48] use RL algorithms to train policies. Nagabandi et al. [47] and Hansen et al. [48] propose to improve the sample efficiency in RL by jointly learning the dynamic model and the control policy. To accelerate the convergence of RL, Rajeswaran et al. [16] introduce demonstration augmented policy gradients (DAPG) and train RL with expert demonstrations. Instead of manually collecting expert demonstrations, ILAD [29] derives demonstration data via an affordance model [49] and trajectory optimization [50]. Xu et al. [38] and Wan et al. [46] decompose the dexterous grasping into two sub-tasks: the grasp proposal generation and the goal-conditioned policy training and largely improve the generalization of the learned policy. However, RL is data-inefficient, requires well-designed rewards, and often results in unrealistic object manipulation. In this work, we leverage human videos to improve RL training of state-based policies and then train a unified visual policy that enables object manipulation given 3D point cloud inputs.
@@
Learning robotic manipulation from human videos. Human videos naturally provide rich human hands interactions with diverse objects [51, 52, 53, 54, 55]. Equipping robots with the capacity to acquire manipulation skills by simply watching human videos has been an attractive direction [56, 57, 58, 59, 60]. One line of works [37, 61, 62, 63, 64] aims to learn generic visual representations from large-scale human video data [65] and then makes use of such pre-trained representations to learn control policies. Another line of works [66, 67, 68] focuses on learning reward functions from human videos. Chen et al. [66] introduce DVD, a domain-agnostic video discriminator, to learn multi-task reward functions. Alakuijala et al. [67] propose a task-agnostic reward function by using unlabeled human videos. Some other works [24, 25, 26, 27, 69] explore how to explicitly leverage human videos to facilitate learning dexterous manipulation skills. Mandikal et al. [25, 69] propose to utilize object affordances priors and human grasping priors to improve the dexterous grasping policy. DexMV [24] and DexRepNet [26] extract demonstrations data from human videos and use DAPG algorithm [16] to launch the RL training with demonstrations. PGDM [28] proposes to initialize the robot hand configuration at the given pre-grasp derived from human motions and enable efficient RL training. Compared to prior work [24, 26], our method allows to solve more complex manipulation tasks with fewer videos by using a unified trajectory guided reward learned from videos and automatic trajectory augmentation.
@@ III Method
This work aims to train vision-based dexterous manipulation policies. However, jointly learning the visual representation and the control policy is challenging. Therefore, we first train the state-based policy and then distill the learned policy to a vision-based policy. To train the state-based policy, as the pure RL training is ineffective [24, 26, 29], we extract reference trajectories from human videos and propose trajectory-guided RL to improve the performance. In the following, we first present the reference trajectory generation in Section III-A. Then, we introduce our state-based policy learning algorithm in Section III-B. Finally, we describe the visual policy and its training in Section III-C.
@@ III-A Reference Trajectory Extraction
To guide our policy training, we extract hand and object poses from video demonstrations. The poses and shapes of hands are represented by MANO [70] and result in 3D locations of joints $\bm{\psi}_{h}\in\mathbb{R}^{21\times 3}$. To account for differences in geometry, we retarget the human hand pose to the robot hand pose. Following [24, 71, 72], we formulate the hand motion retargeting from a video of length $T$ as an optimization problem and define its objective function as:
$\begin{split}\min_{\bm{q}_{r}^{t}}\sum_{t=1}^{T}\Big{\|}\bm{\hat{x}}_{rj}^{t}(\bm{q}_{r}^{t})-\bm{{\psi}}_{hj}^{t}\Big{\|}^{2}_{2}+\alpha\Big{\|}\bm{q}_{r}^{t}-\bm{q}_{r}^{t-1}\Big{\|}^{2}_{2},\end{split}$ (1)
where $\bm{q}_{r}^{t}$ represents robot joint rotation angles. $\bm{{\psi}}_{hj}^{t}$ denotes human hand tip and middle phalanx positions. We solve their robot counterparts $\bm{\hat{x}}_{rj}^{t}$ via forward kinematics according to given ${\bm{q}_{r}^{t}}$. The first term aims to minimize the difference between the 3D locations of robot joints to those in the human hand. We also use a regularization term to avoid sudden changes in the robot pose, where we empirically set $\alpha$ to 4e-3 and ${\bm{q}_{r}^{0}}$ to the mean pose of its motion limits. We use the NLopt solver [73] for the optimization and port robot and object motions into the simulator to generate reference trajectories, which are visually plausible but not physically plausible. Figure 2 illustrates results of the motion retargeting for the Allegro robot hand.
Figure 2: Motion retargeting results for the Allegro hand and objects under different poses for selected DexYCB videos.
@@ III-B State-based Policy
We train a state-based policy with RL to recover physically plausible trajectories, where the above generated reference trajectories are employed in the reward function to guide robot hand and object motions. The network architecture for the state-based policy consists of actor and critic MLPs [74]. It takes both robot and object states as inputs and predicts the robot control commands. In the following, we will describe our reward functions and the trajectory augmentation for RL.
Trajectory-guided reward functions. During RL training, we propose to divide the reference trajectory into two stages: the pre-grasp stage and the manipulation stage.
During the pre-grasp stage, the robot hand needs to approach the object without making physical contacts. We require that the robot approaches the object in a similar way as humans and define the following reward function as:
@@
where $T_{p}$ is the length of pre-grasp steps, and $\bm{\hat{x}}_{rt}^{t}$ denotes the robot finger tip positions at the timestep $t$ in the reference trajectory. $\bm{{x}}_{rt}^{t}$ is the current robot finger tip positions.
@@
When the robot successfully reaches its pre-grasp configuration, the episode starts its manipulation stage. In this stage, the robot aims to manipulate the object and bring it to the desired target configuration. Here, we define the reward function to constrain the robot and object motions jointly:
$\begin{split}R_{m}=\sum_{t=T_{p}+1}^{T_{r}}\lambda_{1}R_{m}^{h}+\lambda_{2}R_{m}^{o}+\lambda_{3}\mathds{1}_{\rm cont}+\lambda_{4}\mathds{1}_{\rm lift},\end{split}$ (3)
@@
where $T_{r}$ is the length of the reference trajectory. The first term $R_{m}^{h}$ constrains the hand motions similar to (2). The second term $R_{m}^{o}={\rm exp}(-\alpha_{1}(\Big{\|}\bm{x}_{o}^{t}-\bm{\hat{x}}_{o}^{t}\Big{\|}^{2}_{2}+\alpha_{2}\phi(\bm{\theta}_{o}^{t},~{}\bm{\hat{\theta}}_{o}^{t})))$ constrains the object motions, where $\bm{x}_{o}^{t}$ and $\bm{\hat{x}}_{o}^{t}$ are the current object position and its reference at timestep $t$. $\phi(\cdot)$ computes the angular distance between the current object orientation with its reference $\bm{\hat{\theta}}_{o}^{t}$. The third term computes the number of finger tips in contact with the object. The forth term assigns bonus points when the object is lifted off the table. We empirically set $\lambda_{1}=4,\lambda_{2}=10,\lambda_{3}=0.5,\alpha_{1}=50$ and $\alpha_{2}=0.1$. Since our reward function is primarily derived from trajectories in videos, it can be applied across various manipulation tasks, reducing the need for task-specific reward engineering as in DexMV [24].
@@
Reference trajectory augmentation. Though the state-based policy can successfully imitate the hand and object motions from a video, it remains challenging to generalize to different initial object positions, rotations and target positions. To make our policy applicable to different initial and target object configurations, we introduce our reference trajectory augmentation strategies during RL training. Specifically, we randomly set initial object positions or rotations and then transform the whole reference trajectory accordingly. In order to further increase the diversity of target object positions, we augment the original object trajectory through interpolations between the original last object pose and the target pose. We similarly interpolate hand motions and use proposed reward functions to train the state-based policy.